### v4

# 1Association for the Advancement of Artificial Intelligence

Abstract
Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky. The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging. In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm. First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem. Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes. Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible. Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation. The code and models will be made publicly available.
Introduction
In fields such as land cover classification [lv2023novel, tong2020land], urban planning [zhu2019understanding, zhao2019exploring], and disaster response [kucharczyk2021remote, zhao2024see, wang2023hierarchical], the analysis of aerial imagery plays a crucial role. The cross-view Ground-to-Aerial Image Synthesis task leverages ground street-view images to synthesize corresponding aerial images at the same location. Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [arrabi2024cross]. Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [ou2023method] or fill historical gaps in high-resolution satellite imagery records [cornebise2022open], among other applications.
Figure 1: Illustration of the cross-view image synthesis task. (a) Challenges of cross-view domain gaps; (b) Challenges of occlusion in dense scenes; (c) Comparing our ground-to-aerial image synthesis method with existing cross-view synthesis methods.
Despite the potential of cross-view synthesis from ground-level to aerial imagery, significant challenges remain. The drastic viewpoint differences between street and satellite views lead to distinct object appearances, resulting in substantial domain gaps across image types. As shown in Figure 1 (a), in cross-view scenarios, street-view images primarily capture ground and building facade details, while satellite views present more rooftop and macroscopic summary information. Even with state-of-the-art diffusion methods, the domain gap between views often results in generated images that are realistic yet not true to reality; the synthesized satellite imagery may differ significantly from the actual content of corresponding street-views, as shown in Figure 1(c).
Another challenge is object occlusion in dense urban settings, where tall buildings or trees limit the visibility range in street-views, which is substantially smaller than the coverage of satellite images, as shown in Figure 1(b). Aerial images synthesized from a single street-view image struggle to maintain consistency with real satellite images in unseen areas. An effective approach is to leverage multiple neighboring street-view images for satellite view synthesis. However, this poses a challenge for the model in effectively utilizing geographic location information from street-views and integrating features from multiple street-view images.
To address these challenges, this paper introduces SkyDiffusion, a ground-to-aerial image synthesis method based on diffusion models and the BEV paradigm, emulating a sky-view perspective. We first propose a Curved-BEV method, which transforms street-view images into Bird’s-Eye view to achieve domain alignment across viewpoints. Additionally, the Curved-BEV method introduces a multi-to-one mapping strategy, effectively integrating street-view images from different capture locations into a unified BEV space, thus overcoming the limited observation range in urban street-view images. Subsequently, we designed a BEV-guided diffusion model to generate realistic satellite images consistent with street-view content. Therefore, SkyDiffusion effectively overcomes the challenges of cross-view domain gaps and the limited observation range in densely occluded urban scenes, generating more realistic satellite images.
Moreover, existing cross-view datasets, such as CVUSA [workman2015wide] and VIGOR [zhu2021vigor], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis. To address this, we introduce the Ground2Aerial-3 (G2A-3) dataset, which includes three novel aerial image generation scenarios: disaster emergency response, low-altitude UAV imagery, and historical high-resolution imagery generation tasks. The Ground2Aerial-3 dataset provides greater practical value and significance for exploring cross-view synthesis tasks in more diverse scenarios. Our main contributions are as follows:
  • We introduce SkyDiffusion, a novel ground-to-aerial synthesis method leveraging diffusion models and BEV paradigm to generate realistic, consistent aerial images.
  • We design a Curved-BEV method to transform street-view images into satellite views for domain alignment. It also includes "Multi-to-One" mapping strategy to enhance BEV perception range in densely occluded urban areas.
  • We introduce Ground2Aerial-3, a new ground-to-aerial image synthesis dataset, featuring disaster scene aerial image, low-altitude UAV image, and historical high-resolution satellite image synthesis tasks.
  • SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), with an SSIM increase of 8.67% and a FID reduction of 21.50%.
Related Work
Cross-view image synthesis involves generating realistic scene images with significant viewpoint changes [shi2022geometry, lu2020geometry, Sat2Density, li2024crossviewdiff, xu2024geospecific]. This task is common in computer vision and has broad applications in areas such as remote sensing image processing, virtual reality, and autonomous driving. It is mainly divided into aerial-to-ground and ground-to-aerial generation [X-Seq, regmi2019cross]. In this field, synthesizing satellite images from street-views is a challenging yet important task. X-Seq attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies. tang2019multi used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios. Additionally, the performance of GANs has also limited the quality of ground-to-aerial image synthesis in previous studies. Concurrent work [arrabi2024cross] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images. Unlike our method, it uses existing BEV methods and only supports one-to-one synthesis.
Figure 2: Overview of the proposed SkyDiffusion framework. It include the curved BEV transformation and BEV-controlled diffusion model. The lower parts present the results of One-to-One and Multi-to-One BEV transformations, respectively.
BEV transformation is a perspective domain conversion technique similar to polar transformation [shi2019spatial, toker2021coming] and geo-transformation [lu2020geometry], widely used in retrieval [ye2024cross], navigation [bevformer] and localization tasks [song2024learning, wang2024fine, shi2023boosting]. Traditional BEV transformation techniques, which rely on camera parameters [sarlin2023orienternet] or depth estimation methods [teng2024360bev, ye2024sg], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation. Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [wang2024fine] or feature level [shi2023boosting]. However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss. Moreover, most existing panoramic BEV transformation methods primarily focus on one-to-one mappings, with relatively little research on multi-to-one BEV joint mappings.
Novel view synthesis based on diffusion models. Diffusion models have shown significant potential in recent studies on novel view synthesis [kawar2023imagic, gao2023magicdrive, consis]. DreamBooth [ruiz2023dreambooth] and AerialDiffusion [AerialDiffusion] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models. However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone. Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [li2023bbdm, insd, control]. However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
Methods
As illustrated in Fig. 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images. It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section Curved BEV Transformation). Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images (Section BEV-Controlled Diffusion Model). Following this, we detail our network training process (Section Network Training and Implementation Details).
Curved BEV Transformation
One-to-One BEV. Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[shi2023boosting, wang2024fine]. However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area. Thus, we propose an enhanced curved BEV transformation to better capture cross-view information and address this issue.
As shown in Fig. 3, we use $P(i,j)$ to denote an index on the BEV space, and $P(x,y,z)$ to illustrate a point in a three-dimensional coordinate system with the camera on the z-axis. We assume that the BEV plane is an upward curved surface, with the height (z-axis) increasing rapidly as the 3D points move away from the center (the central part of Fig 3(a)). The following equation formulates the relationship between $P(x,y,z)$ and $P(i,j)$, where $d_{norm}$ represents the normalized distance from the center, $d_{max}$ is the maximum distance from the center, and $\lambda$ is a scaling factor.
$\begin{cases}x=j-\frac{l}{2},\quad y=\frac{l}{2}-i\\ z=d_{norm}^{4}\times\lambda=\left(\frac{\sqrt{x^{2}+y^{2}}}{d_{max}}\right)^{4}\times\lambda\end{cases}$ (1)
Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$. Based on the equirectangular projection properties of street-view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$. This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image. More details are in the supplementary material.
{$u&=\left[\text{arctan2}\left(y,x\right)+\pi\right]\frac{w}{2\pi}\\ v&=\left[\frac{\pi}{2}+\text{arctan2}\left(z-H,\sqrt{x^{2}+y^{2}}\right)\right]\frac{h}{\pi}\end{aligned}\right.$ (2)
Following Eq. 1 and Eq. 2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [teng2024360bev, ye2024sg]. This effectively maps upper street-view features like distant roads and buildings. Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
Figure 3: Schematic of the curved BEV transformation. It illustrates the mapping of two points on the BEV plane to the top and bottom of the street-view image during transformation.
Multi-to-One BEV. Due to the limitations imposed by dense buildings on street-view observations, the BEV sensing range is very limited in urban datasets like VIGOR. To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street-view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position. We select the street-view mapping result closest to the capture point for overlapping areas following Eq. 3.
$&\textit{BEV}\left(x,y\right)=\textit{BEV}_{\mathrm{cam}_{k}}\left(x-\Delta x_{k},y-\Delta y_{k}\right)\\ &k=\arg\min_{i}\sqrt{(x-x_{\mathrm{cam}_{i}})^{2}+(y-y_{\mathrm{cam}_{i}})^{2}}\end{aligned}\right.$ (3)
where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system. Additionally, during the implementation process, we also include boundary cropping as part of the image post-processing.
BEV-Controlled Diffusion Model
Image diffusion models progressively denoise images and generate samples from the training domain. The denoising process can occur in pixel space or in a latent space encoded from training data [ddim, ddpm]. In the forward process, diffusion models gradually add Gaussian noises to a ground truth image $x_{0}$ according to a predetermined schedule $\beta_{1},\beta_{2},\dots,\beta_{T}$:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (4)
where $x_{t}$ is a noised sample with noise level $t$. The reverse process involves a series of denoising steps, where noise is progressively removed by employing a neural network $\epsilon_{\theta}$ with parameters $\theta$. The reverse process involves a series of denoising steps, where noise is progressively removed by employing a neural network $\epsilon_{\theta}$ with parameters $\theta$. This neural network predicts the noise $\epsilon$ present in a noisy image $x_{t}$ at step $t$.
In the satellite image synthesis process, an effective feature embedding mechanism is required to ensure the generated satellite image aligns with the BEV image $I_{bev}$. We use a conditional encoder to transform the BEV image into a latent space feature embedding, and then, similar to ControlNet [control], inject the features into the diffusion model using zero convolution layers and replicated SD modules. Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$. Notably, considering the potential feature distortions in BEV images, we employ a spatial attention mechanism to enhance feature capture while suppressing distorted features.
$c_{bev}=\mathcal{E}(I_{bev}).$ (5)
Subsequently, by using zero convolution layers and replicating the encoder and middle block structures and weights from the SD module, the extracted BEV features are injected into another SD module. Guided by the BEV image, the diffusion model incorporates cross-view associative information, synthesizing aerial views with consistent content.
Network Training and Implementation Details
Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image. It then learns a network $\epsilon_{\theta}$ to predict the noise added to the noisy satellite image $Z_{t}$ based on the given BEV image as the conditional input. Since text cannot fully describe the street-view scene, our cross-view image synthesis is essentially an image-conditioned synthesis task, with text not being our optimization target. This helps the model focus on generating satellite images based on the BEV conditions.
$\mathcal{L}=\mathbb{E}_{sat_{0},t,c_{bev},\epsilon\sim\mathcal{N}(0,1)}\left[\left\|\epsilon-\epsilon_{\theta}(sat_{t},t,c_{bev})\right\|_{2}^{2}\right]$ (6)
The loss $\mathcal{L}$ is the learning objective of the diffusion model, where $t$ represents the time step. We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [stable], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [cfg]. Final inference sampling uses 50 steps with the DDIM strategy [ddim]. Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs.